# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v2

@@ 4.1 Vanilla Training
This classifier’s predicted label for a sample $\bm{x}\in\mathbb{R}^{d}$ is ${\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}_{0}\rangle)$. Note that $\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle=(1/n)\sum_{i}\widehat{y}_{i}\langle\bm{x}_{i},\bm{x}\rangle$ is an average of noisy labels in the training set with weights given by $\langle\bm{x}_{i},\bm{x}\rangle/n$. Therefore, it is similar to kernel methods with the inner product kernel.
Our next result bounds the probability of ${\widehat{\bm{\theta}}}_{0}$ correctly classifying a given fixed test point $\bm{x}$.
Theorem 4.1 (Vanilla Training)
@@
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)=y)$ $\leq\alpha_{0}(\bm{x}):=1-\frac{1}{2\sqrt{2\pi}}\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\,\text{ and }$
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)=y)$ $\geq\widetilde{\alpha}_{0}(\bm{x}):=1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
@@
[…]
We will next define the accuracy of a classifier ${\widehat{\bm{\theta}}}$ as
${\sf acc}({\widehat{\bm{\theta}}}):=\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}\rangle)=y)\,,$ (4.4)
where the probability is with respect to the randomness in the training set used to learn ${\widehat{\bm{\theta}}}$ as well as $(\bm{x},y)$. Our next result bounds the accuracy of the classifier ${\widehat{\bm{\theta}}}_{0}$ obtained by vanilla training.
Theorem 4.2 (Vanilla Training: Population Accuracy)
@@
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\leq 1-\frac{1}{4\sqrt{2\pi}}\exp\left(-160(1+\sqrt{n}(1-2p))^{2}\frac{\gamma^{4}}{\lambda_{\min}^{2}d}\right)\left(1-e^{-d/16}\right)\,\text{ and }$ (4.5)
${\sf acc}({\widehat{\bm{\theta}}}_{0})$ $\geq 1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\gamma^{4}}{16\lambda_{\max}^{2}d}\right)-e^{-d/8}-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,,$ (4.6)
for an absolute constant $c>0$.
@@
Remark 4.3 (Tightness of Accuracy Bounds)
When $n\lesssim\frac{d^{2}}{(1-2p)^{2}\gamma^{4}}$, $\gamma\lesssim d^{1/4}$ and $\frac{\lambda_{\max}}{\lambda_{\min}}\lesssim 1$, then the lower and upper bounds for $\textup{acc}(\widehat{\bm{\theta}}_{0})$ in Theorem 4.2 match (up to constant factors).
@@
Corollary 4.4 (Vanilla Training: Sample Complexity)
For any $\delta>2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$, having sample size $n\geq 8\lambda_{\max}^{2}\frac{\log{1}/{\delta}}{(1-2p)^{2}}\frac{d}{\gamma^{4}}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-\delta$. In particular, when the label flipping probability satisfies $p>2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$, then $n\geq 8\lambda_{\max}^{2}\frac{\log{1}/{p}}{(1-2p)^{2}}\frac{d}{\gamma^{4}}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-p$, i.e., our learned classifier $\widehat{\bm{\theta}}_{0}$ has better accuracy than the source providing noisy labels (used to learn $\widehat{\bm{\theta}}_{0}$).
Remark 4.5 (Effect of Degree of Separation)
As the parameter quantifying the degree of separation $\gamma$ decreases, the accuracy bound in Theorem 4.2 also decreases and the sample complexity required to outperform the noisy label source in Corollary 4.4 increases. This is consistent with our intuition that a classification task should become harder as the degree of separation reduces; we also saw this in Figure 1.
@@
Theorem 4.6 (Information-Theoretic Lower Bound on Sample Complexity)
With a slight generalization of notation, let ${\sf acc}({\widehat{\bm{\theta}}};{\bm{{\mu}}})$ denote the accuracy of the classifier ${\widehat{\bm{\theta}}}$, when the ground truth separator is ${\bm{{\mu}}}$. We also consider the case of $\bm{\Sigma}=\mathcal{P}^{\perp}_{\bm{\mu}}$, viz., the projection matrix onto the space orthogonal to ${\bm{{\mu}}}$. For any classifier ${\widehat{\bm{\theta}}}$ learned from $\mathcal{T}:=\{(\bm{x}_{j},\widehat{y}_{j})\}_{j\in[n]}$, in order to achieve $\inf_{{\bm{{\mu}}}\in\mathbb{S}^{d-1}}{\sf acc}({\widehat{\bm{\theta}}};{\bm{{\mu}}})\geq 1-\delta$, the condition $n=\Omega\left(\frac{(1-\delta)}{(1-2p)^{2}}d\right)$ is necessary in our problem setting.
It is worth mentioning that there is a similar lower bound in Gentile and Helmbold, (1998) for a different classification setting. In contrast, Theorem 4.6 is tailored to our setting and moreover, the proof technique is also different and interesting in its own right. Specifically, for the proof of Theorem 4.6, we follow a standard technique in proving minimax lower bounds which is to reduce the problem of interest to an appropriate multi-way hypothesis testing problem; this is accompanied by the application of the conditional version of Fano’s inequality and some ideas from high-dimensional geometry. This proof is in Appendix D.
Remark 4.7 (Minimax Optimality of Sample Complexity)
@@ 4.2 Retraining
We first label the training set using ${\widehat{\bm{\theta}}}_{0}$. Denote by $\widetilde{y}_{i}={\rm sign}(\langle\bm{x}_{i},{\widehat{\bm{\theta}}}_{0}\rangle)$ the predicted label for sample $\bm{x}_{i}$. We then retrain the model using these predicted labels. Our learned classifier here is similar to the one in Section 4.1, except that the observed labels are replaced by the predicted labels. Specifically, our retraining classifier is the following:
@@
Our retraining classifier’s predicted label for a sample $\bm{x}\in\mathbb{R}^{d}$ is ${\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}_{1}\rangle)$. Our next result bounds the probability of ${\widehat{\bm{\theta}}}_{1}$ correctly classifying a given fixed test point $\bm{x}$.
Theorem 4.8 (Retraining)
@@
Suppose that $\frac{n}{d}>\frac{4\lambda_{\max}}{\gamma^{2}(1-2p)}$ and $nd>\frac{\gamma^{4}}{\lambda_{\max}^{2}}$. Consider $\bm{x}\notin\mathcal{T}$ and let $y$ be its true label. We have
@@
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle)=y)\geq\alpha_{1}(\bm{x})\,,$
where
@@
$\alpha_{1}(\bm{x}):=1$ $-2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$-4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)-\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,,$
with
$q^{\prime}=\exp\left(-\frac{n(1-2p)\gamma^{2}}{40\lambda_{\max}}\right)\text{ and }p^{\prime}=\left(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd}\right)p\,.$
The proof of Theorem 4.8 is in Appendix E.
@@
Technical Challenges. The proof of Theorem 4.8 is especially challenging because as we discussed after Theorem 4.1, the classifier learned with vanilla training, i.e., ${\widehat{\bm{\theta}}}_{0}$, is a non-uniform noisy label provider. In addition, ${\widehat{\bm{\theta}}}_{0}$ depends on all the samples in the training set and hence each predicted label $\widetilde{y}_{i}$ is also dependent on the entire training set. This dependence of the predicted labels on all the data points makes the analysis even more challenging because standard arguments with independence are not applicable here. In contrast, the analysis of ${\widehat{\bm{\theta}}}_{0}$ was technically simpler as it was based on noisy labels $\widehat{y}_{i}$’s that were independent across samples. The high-level proof idea for Theorem 4.8 is that for each sample $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$ in the training set, we carefully curate a dummy label $\widetilde{y}^{\prime}_{\ell}$ which does not depend on ${\bm{z}}_{k}$ for $k\neq\ell$, but depends on all $\{u_{k}\}_{k\in[n]}$; see (E.2) in Appendix E. We show that with high probability the labels predicted by the vanilla estimator ${\widehat{\bm{\theta}}}_{0}$ on the training samples, i.e. ${\rm sign}(\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle)$ are the same as the dummy labels $\widetilde{y}^{\prime}_{\ell}$. Next, under this high probability event, we use the dummy labels instead, which are more amenable to analysis as they are independent of $\{{\bm{z}}_{k}\}_{k\neq\ell}$. At the beginning of Appendix E, we provide an outline of these steps in more detail.
It is worth noting that the lower bound derived on $\alpha_{1}(\bm{x})$ is increasing in $\langle\bm{x},{\bm{{\mu}}}\rangle$. So, just like the vanilla training classifier ${\widehat{\bm{\theta}}}_{0}$, the classifier ${\widehat{\bm{\theta}}}_{1}$ learned with retraining is more likely to be wrong on the samples that are less aligned with the ground truth separator ${\bm{{\mu}}}$. We will now provide a lower bound on the accuracy (defined in (4.4)) of the classifier ${\widehat{\bm{\theta}}}_{1}$.
Theorem 4.9 (Retraining: Population Accuracy)
Under the setting of Theorem 4.8 we have
@@
${\sf acc}({\widehat{\bm{\theta}}}_{1})\geq 1$ $-2\exp\left(-\frac{n(1-2q^{\prime})^{2}\gamma^{4}}{64(\gamma^{4}+2\lambda_{\max}^{2}d)}\right)-2e^{-d/8}$
$-4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)-\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$ (4.8)
@@
The proof of Theorem 4.9 is in Appendix F. As expected, the lower bound (4.8) is increasing in $\gamma$, because as the degree of separation increases, the classification task becomes easier.
@@
Accuracy of Retraining vs. Vanilla Training. We will now compare the accuracy of vanilla training model ${\widehat{\bm{\theta}}}_{0}$ with that of the retrained model ${\widehat{\bm{\theta}}}_{1}$, identifying regimes where retraining improves the model accuracy. Before proceeding further, it is useful to note that the dominant term in $1-{\sf acc}({\widehat{\bm{\theta}}}_{1})$ is the last term in (4.8) which has $(1-2p^{\prime})$ inside the exponent and $p^{\prime}$ can become arbitrarily close to $p$ as $n$ or $d$ grows. In contrast, the upper bound on ${\sf acc}({\widehat{\bm{\theta}}}_{0})$ in (4.5) has $(1-2p)^{2}$ inside the exponent. This is the main observation which indicates that retraining can improve model accuracy, and this improvement becomes increasingly significant as $p$ approaches $\frac{1}{2}$. We will now formalize this observation by providing sufficient conditions for this improvement to happen.
Remark 4.10 (When Does Retraining Improve Accuracy?)
@@
We consider an asymptotic regime where $n,d\to\infty$. Also, suppose that $\lambda_{\min}$ and $\lambda_{\max}$ remain bounded from above and below by positive constants. We would like to characterize a regime where $1-{\sf acc}({\widehat{\bm{\theta}}}_{1})\leq 1-{\sf acc}({\widehat{\bm{\theta}}}_{0})$, i.e., under what conditions the misclassification error of ${\widehat{\bm{\theta}}}_{1}$ is smaller than that of ${\widehat{\bm{\theta}}}_{0}$. By comparing the bounds in Theorems 4.2 and 4.9, we obtain the following sufficient conditions. There exists an absolute constant $c_{0}$ such that if $p\in(\frac{1}{2}-c_{0},\frac{1}{2})$ and
$\frac{d}{\gamma^{2}(1-2p)}\max\left\{1,\frac{\log n}{\gamma^{2}}\right\}\lesssim n\lesssim\frac{d^{2}}{\gamma^{4}(1-2p)^{2}}\,,$ (4.9)
then the lower bound on the accuracy of retraining (in eq. 4.8) is greater than the upper bound on the accuracy of vanilla training (in eq. 4.5).
Note that as the degree of separation $\gamma$ decreases, the range of $n$ in (4.9) becomes larger but also shifts up, indicating we need more samples (keeping other parameters fixed) to obtain an improvement with retraining.

### v3

@@ 4.1 Initial Training
This classifier’s predicted label for a sample $\bm{x}\in\mathbb{R}^{d}$ is ${\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}_{0}\rangle)$. Note that $\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle=\frac{1}{n}\sum_{i}\widehat{y}_{i}\langle\bm{x}_{i},\bm{x}\rangle$ is a weighted average of the noisy labels in the training set, with the weights being $\frac{1}{n}\langle\bm{x}_{i},\bm{x}\rangle$. So it is similar to kernel methods with the inner product kernel.
Our next result bounds the probability of ${\widehat{\bm{\theta}}}_{0}$ misclassifying a fixed test point $\bm{x}$, with respect to the randomness in the training set $\mathcal{T}$.
Theorem 4.1 (Initial training).
@@
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)\geq\alpha_{0}(\bm{x}):=$
$\frac{1}{2\sqrt{2\pi}}\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\,\text{ and }$
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)\leq\widetilde{\alpha}_{0}(\bm{x}):=$
$\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)+2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
@@
$\widetilde{y}=\begin{cases}y\text{ with probability at least }\geq 1-\widetilde{\alpha}_{0}(\bm{x})\,\text{ and}\\ -y\text{ with probability at most }\leq\widetilde{\alpha}_{0}(\bm{x}),\end{cases}$
where $\widetilde{\alpha}_{0}(\bm{x})$ is as defined in Theorem 4.1. We will next define the population error of a classifier ${\widehat{\bm{\theta}}}$ as
${\sf err}({\widehat{\bm{\theta}}}):=\operatorname{\mathbb{E}}_{(\bm{x},y)}\big{[}\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}\rangle)\neq y)\big{]},$ (4.4)
where the probability (inside the expectation) is w.r.t. the randomness in the training set used to learn ${\widehat{\bm{\theta}}}$. Similarly, the population accuracy of ${\widehat{\bm{\theta}}}$ is defined as
${\sf acc}({\widehat{\bm{\theta}}}):=\operatorname{\mathbb{E}}_{(\bm{x},y)}\big{[}\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}\rangle)=y)\big{]}=1-{\sf err}({\widehat{\bm{\theta}}}).$ (4.5)
Depending on the context, our results and discussions will sometimes be presented in terms of error (lower is better) and at other times in terms of accuracy (higher is better).
Our next result bounds the error of the classifier ${\widehat{\bm{\theta}}}_{0}$.
Theorem 4.2 (Initial training’s population error).
@@
${\sf err}({\widehat{\bm{\theta}}}_{0})\geq\frac{\left(1-e^{-\frac{d}{16}}\right)}{4\sqrt{2\pi}}\exp\left(-\frac{160(1+\sqrt{n}(1-2p))^{2}\gamma^{4}}{\lambda_{\min}^{2}d}\right),$ (4.6)
${\sf err}({\widehat{\bm{\theta}}}_{0})\leq\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\gamma^{4}}{16\lambda_{\max}^{2}d}\right)+e^{-\frac{d}{8}}\\ +2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (4.7)
@@
Remark 4.3 (Tightness of error bounds).
When $n\lesssim\frac{\lambda_{\max}^{2}d^{2}}{(1-2p)^{2}\gamma^{4}}$, $\gamma\lesssim(\lambda_{\max}^{2}d)^{1/4}$ and $\frac{\lambda_{\max}}{\lambda_{\min}}\lesssim 1$, then the lower and upper bounds for ${\sf err}({\widehat{\bm{\theta}}}_{0})$ in Theorem 4.2 match (up to constant factors).
@@
Corollary 4.4 (Initial training’s sample complexity).
For any $\delta>2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$, having number of samples $n\geq 8\lambda_{\max}^{2}\frac{\log{1}/{\delta}}{(1-2p)^{2}}\frac{d}{\gamma^{4}}$ ensures ${\sf acc}(\widehat{\bm{\theta}}_{0})>1-\delta$. In particular, when the label flipping probability satisfies $p>2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$, then $n\geq 8\lambda_{\max}^{2}\frac{\log{1}/{p}}{(1-2p)^{2}}\frac{d}{\gamma^{4}}$ ensures $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-p$, i.e., our learned classifier $\widehat{\bm{\theta}}_{0}$ has better accuracy than the source providing noisy labels (used to learn $\widehat{\bm{\theta}}_{0}$).
Remark 4.5 (Effect of degree of separation).
As the parameter quantifying the degree of separation $\gamma$ decreases, the error bound in Theorem 4.2 increases and the sample complexity required to outperform the noisy label source in Corollary 4.4 increases. This is consistent with our intuition that a classification task should become harder as the degree of separation reduces; we saw this in Figure 1.
@@
Theorem 4.6 (Information-theoretic lower bound on sample complexity).
With a slight generalization of notation, let ${\sf acc}({\widehat{\bm{\theta}}};{\bm{{\mu}}})$ denote the accuracy of the classifier ${\widehat{\bm{\theta}}}$ as per Equation 4.5, when the ground truth separator is ${\bm{{\mu}}}$. We also consider the case of $\bm{\Sigma}=\mathcal{P}^{\perp}_{\bm{\mu}}$, viz., the projection matrix onto the space orthogonal to ${\bm{{\mu}}}$. For any classifier ${\widehat{\bm{\theta}}}$ learned from $\mathcal{T}:=\{(\bm{x}_{j},\widehat{y}_{j})\}_{j\in[n]}$, in order to achieve $\inf_{{\bm{{\mu}}}\in\mathbb{S}^{d-1}}{\sf acc}({\widehat{\bm{\theta}}};{\bm{{\mu}}})\geq 1-\delta$, the condition $n=\Omega\left(\frac{(1-\delta)}{(1-2p)^{2}}d\right)$ is necessary in our problem setting.
It is worth mentioning that there is a similar lower bound in Gentile & Helmbold (1998) for a different classification setting. In contrast, Theorem 4.6 is tailored to our setting and moreover, the proof technique is also different and interesting in its own right. Specifically, for the proof of Theorem 4.6, we follow a standard technique in proving minimax lower bounds which is to reduce the problem of interest to an appropriate multi-way hypothesis testing problem; this is accompanied by the application of the conditional version of Fano’s inequality and some ideas from high-dimensional geometry. This proof is in Appendix D. There are also a few other noteworthy lower bounds (Cai & Wei, 2021; Im & Grigas, 2023; Zhu et al., 2024) in more general settings than ours; naturally, their bounds are weaker than Theorem 4.6. We discuss these bounds in Appendix E.
Remark 4.7 (Minimax optimality of sample complexity).
@@ 4.2 Retraining
We first label the training set using ${\widehat{\bm{\theta}}}_{0}$. Denote by $\widetilde{y}_{i}={\rm sign}(\langle\bm{x}_{i},{\widehat{\bm{\theta}}}_{0}\rangle)$ the predicted label for sample $\bm{x}_{i}$. We then retrain using these predicted labels. Our learned classifier here is similar to the one in Section 4.1, except that the observed labels are replaced by the predicted labels. Specifically, our retraining classifier model is the following:
@@
${\widehat{\bm{\theta}}}_{1}$’s predicted label for a sample $\bm{x}\in\mathbb{R}^{d}$ is ${\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}_{1}\rangle)$. Our next result bounds the probability of ${\widehat{\bm{\theta}}}_{1}$ misclassifying a fixed test point $\bm{x}$, with respect to the randomness in the training set $\mathcal{T}$ (similar to Theorem 4.1).
Theorem 4.8 (Retraining).
@@
Suppose that $\frac{n}{d}>\frac{4\lambda_{\max}}{\gamma^{2}(1-2p)}$, $nd>\frac{\gamma^{4}}{\lambda_{\max}^{2}}$, and $d\geq 7$. Consider $\bm{x}\notin\mathcal{T}$ and let $y$ be its true label. We have
@@
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle)\neq y)\leq\alpha_{1}(\bm{x}),\text{ where}$
@@
$\alpha_{1}(\bm{x})$ $:=2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$
$+\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,,$
with $q^{\prime}=\exp\left(-\frac{n(1-2p)\gamma^{2}}{40\lambda_{\max}}\right)$ and $p^{\prime}=\left(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd}\right)p$.
The proof of Theorem 4.8 is in Appendix F.
Technical challenges: The proof of Theorem 4.8 is especially challenging due to the following reasons:
(i) The source of the predicted labels $\{\widetilde{y}_{i}\}_{i=1}^{n}$ used by ${\widehat{\bm{\theta}}}_{1}$, viz., ${\widehat{\bm{\theta}}}_{0}$, is a non-uniform noisy label provider, as we discussed after Theorem 4.1.
(ii) More importantly, ${\widehat{\bm{\theta}}}_{0}$ depends on all the samples in the training set and hence each predicted label $\widetilde{y}_{i}$ is dependent on the entire training set. This dependence of the predicted labels on all the data points makes the analysis highly challenging because standard techniques under independence are not applicable here. [^6] The analysis of ${\widehat{\bm{\theta}}}_{0}$ (Theorem 4.1) is relatively simpler because the given noisy labels $\{\widehat{y}_{i}\}_{i=1}^{n}$ are independent across samples.
@@
The high-level proof idea for Theorem 4.8 is that for each sample $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$ in the training set, we carefully curate a dummy label $\widetilde{y}^{\prime}_{\ell}$ that does not depend on $\{{\bm{z}}_{k}\}_{k\neq\ell}$ but depends on all $\{u_{k}\}_{k\in[n]}$; see (F.3) in Appendix F. We show that with high probability, the labels predicted by ${\widehat{\bm{\theta}}}_{0}$ on the training set $\{\widetilde{y}_{\ell}\}_{\ell\in[n]}$ match the corresponding dummy labels $\{\widetilde{y}^{\prime}_{\ell}\}_{\ell\in[n]}$. We analyze the misclassification error of ${\widehat{\bm{\theta}}}_{1}$ under this high probability “good” event with the dummy labels $\{\widetilde{y}^{\prime}_{\ell}\}_{\ell\in[n]}$, which are more conducive to analysis (because they do not depend on $\{{\bm{z}}_{k}\}_{k\neq\ell}$), and also bound the probability of the “bad” event (i.e., $\exists$ at least one $\ell$ such that $\widetilde{y}_{\ell}\neq\widetilde{y}^{\prime}_{\ell}$). We provide an outline of these steps in detail at the beginning of Appendix F.
We will now provide an upper bound on the error (defined in (4.4)) of ${\widehat{\bm{\theta}}}_{1}$.
Theorem 4.9 (Retraining’s population error).
Suppose the conditions of Theorem 4.8 hold. Then, we have
@@
${\sf err}({\widehat{\bm{\theta}}}_{1})\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\gamma^{4}}{64(\gamma^{4}+2\lambda_{\max}^{2}d)}\right)\\ +2e^{-d/8}+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\\ +\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n},$ (4.9)
with $q^{\prime}=\exp\left(-\frac{n(1-2p)\gamma^{2}}{40\lambda_{\max}}\right)$ and $p^{\prime}=\left(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd}\right)p$.
@@
The proof of Theorem 4.9 is in Appendix G. As expected, (4.9) is decreasing in $\gamma$, because the classification task becomes easier as the degree of separation increases.
@@
Retraining vs. Initial training. We will now compare the error bounds of the initial training model ${\widehat{\bm{\theta}}}_{0}$ (4.6) from Theorem 4.2 with that of the retraining model ${\widehat{\bm{\theta}}}_{1}$ (4.9) from Theorem 4.9. Note that the dominant term in the upper bound on ${\sf err}({\widehat{\bm{\theta}}}_{1})$ is the last term in (4.9) which has $(1-2p^{\prime})$ inside the exponent, and $p^{\prime}$ approaches $p$ as $n$ and $d$ grow. In contrast, the lower bound on ${\sf err}({\widehat{\bm{\theta}}}_{0})$ in (4.6) has $(1-2p)^{2}$ inside the exponent. This is the main observation that indicates retraining can improve model accuracy, and this improvement becomes increasingly significant as $p\to\frac{1}{2}$. We will now formalize this observation by providing sufficient conditions under which ${\sf err}({\widehat{\bm{\theta}}}_{1})<{\sf err}({\widehat{\bm{\theta}}}_{0})$ or equivalently, ${\sf acc}({\widehat{\bm{\theta}}}_{1})>{\sf acc}({\widehat{\bm{\theta}}}_{0})$.
Remark 4.10 (When does retraining improve accuracy?).
@@
We consider an asymptotic regime where $n,d\to\infty$. Also, suppose that $\frac{\lambda_{\max}}{\lambda_{\min}}\lesssim 1$. By comparing (4.6) with (4.9), we obtain the following sufficient conditions. There exists an absolute constant $c_{0}$ such that if $p\in(\frac{1}{2}-c_{0},\frac{1}{2})$ and
$\frac{\lambda_{\min}^{2}d}{\gamma^{4}(1-2p)^{2}}\log\left(\frac{\lambda_{\min}^{2}d}{\gamma^{4}(1-2p)^{2}}\right)\lesssim n\lesssim\frac{\lambda_{\min}^{2}d^{2}}{\gamma^{4}(1-2p)^{2}},$
then the accuracy of retraining is greater than the accuracy of initial training.
Some comments regarding Remark 4.10 are in order.
Range of $n$ in Remark 4.10. We believe that the lower bound on $n$ is tight w.r.t. $d$ and $p$, ignoring log factors. We claim this because $\Omega\left(\frac{d}{(1-2p)^{2}}\right)$ samples are necessary for $\textup{acc}({\widehat{\bm{\theta}}}_{0})>1-p$ as per Theorem 4.6, and we can only expect the retraining classifier ${\widehat{\bm{\theta}}}_{1}$ to be better than the initial training classifier ${\widehat{\bm{\theta}}}_{0}$ if the accuracy of ${\widehat{\bm{\theta}}}_{0}$’s predicted labels (which is used to train ${\widehat{\bm{\theta}}}_{1}$) is more than the accuracy of the given labels $=1-p$. In contrast, the upper bound on $n$ may be an artifact of our analysis. Having said that, the range of $\frac{d^{2}}{(1-2p)^{2}}\lesssim n$ is not very interesting because it is far more than the number of learnable parameters $=d$.
Effect of degree of separation $\gamma$. As $\gamma$ increases, the minimum value of $n$ in Remark 4.10 for which we obtain an improvement with retraining reduces. This is consistent with the high-level insight of Fig. 1, viz., retraining is more beneficial when the separation between the classes is large.